OpenStudy (anonymous):

Suppose that f(x) varies directly with x and f(x) = 160 when x = 5. What is f(x) when x = 0.5?

OpenStudy (ranga):

f(x) varies directly with x means: f(x) = kx where k is a constant. when x = 5, f(x) = 160. Plug it into f(x) = kx. 160 = k(5). So k = 160/5 = 32 So f(x) = 32x What is f(x) when x = 0.5 f(x) = 32(0.5) = 16.

OpenStudy (ranga):

A simpler way to do it is by proportion. x = 0.5 is one tenth of x = 5. Since this is a direct proportion problem, f(x) at x = 0.5 must also be one tenth of what f(x) is at x = 5 which is 1/10 * 160 = 16.